solve by completing the square: x^2+7/2x=7
Answer by edjones(8007)

x^2+(7/2)x=7 I assume that this is the problem.
x^2+(7/2)x+49/16=7 + 49/16
(4x+7)^2/16=161/16
(4x+7)/4=+-sqrt(161)/4
4x+7=+-sqrt(161)
4x=-7+-sqrt(161)
x=(-7+-sqrt(161))/4
.
Ed